8-pack control system working group meeting

7 January 2003

In attendence: Chestnut, Kim, Nelson, Rock, S. Smith

Notes by: J Nelson

 

Generic agenda was status of the following:

 

Info from 10 a.m. meeting

PLC Checkout

LLRF SW module support

LLRF databases

Other progress this past week

 

Janice gave a brief summary of 8-pack status: 3rd klystron is here, 4th gets installed this week, hope to do more LLRF checkout the end of next week.

 

Ron asked about converting ADC counts to powers. Steve gave the following formula for PWR variables:

Power = (ADCn-PEDn)*GAINn

Kukhee is working on the ADC database. Its getting close. He wishes to do some invasive investigating in NLCTA IOCs. Janice said shed email him when NLCTA is down later this week or next.

 

Janice noticed some of the SIS waveform variables dont seem to exist - have white boxes on the plots. The plot pack_dm2k_APrdbk_calib.adl is one example.

 

Steve confirmed that blind control exists and is fine for the checkout scheduled for the end of next week.

 

Janice will ask Keith who will do the klystron run up for 8-pack do we want an epics algorithm, etc.

 

DIO still on hold. Sooner rather than later Steve and Andy will want to be able to switch the multiplexer for the SISes. Judy will look into existing documentation and existing database templates. Janice and Steve will work with Andy to get what multiplexers will do what and design panels.

(Janice has in the notes from the 12/10/02 meeting that there are buttons on the IQA module panels. Are these related?)

 

Steve will also ask Andy about how the DIO is set up (voltages of inputs, etc)

 

Janice will talk to Joe Frisch about what we want the DIO channel for turning off the modulator to do.

 

Janice will check the monitor on the dance floor to see if its busted.

 

Steve reported that Elias wired the VSTEP modules backwards: 0-15 is really 15-0. Kukhee will fix the database, and Steve will share some of his coffee won in a bet with Kukhee for the inconvenience.

 

Steve and Janice met with Tonee after the meeting. Tonee thinks it will be reasonably easy to ship the waveforms to EPICS. She and Janice will look into setting the box up in the hall. Janice will work with Ron to make sure the database exists for the peak power meter variables.

 

Next week:

ADC progress

DIO progress

Peak Power Meter progress

Other progress reports and issues